A 12 liters mixture is 60% alcohol. If 8 liters of a 25% alcohol mixture are added what percent mixture will be formed
Answer by stanbon(75887)

A 12 liters mixture is 60% alcohol. If 8 liters of a 25% alcohol mixture are added what percent mixture will be formed
-----
Equation:
alcohol + alcohol = alcohol
0.60*12 + 0.25*8 = x*20
---
60*12 + 25*8 = x*2000
-----
720 + 200 = 2000x
----
x = 920/2000
---
x = 0.46 = 46%
===================
Cheers,
Stan H.
===================